# Estimating Detector Error Models on Google’s Willow

### v1

@@ I Introduction
Over the last year, teams have published results directly demonstrating improved performance of logically encoded circuits reichardt2024demonstrationquantumcomputationerror and increased lifetimes of error corrected states acharya2024quantumerrorcorrectionsurface. This leads naturally to an increased interest in the analysis of the resulting syndrome data and the description of the data generating process via detector error models (DEMs).
@@
We then provide algorithms for two learning tasks: a) rate estimation, in which the structure of the DEM is given and the goal is to learn the parameters from syndrome data-sets, and b) structure learning, where both structure and parameters must be learned. For each task, we describe two algorithmic approaches, one based on moments of detectors and the other based on parities of the same. In the last section, we demonstrate the application of these algorithms to recent Google experiments acharya2024quantumerrorcorrectionsurface and on simulated data. Specifically, we discuss the accuracy of decoding using estimated priors versus non-informative priors and priors trained to minimize the logical error rate (LER). Then, we evaluate the agreement between various DEMs and syndromes using a method for estimating the likelihood of syndromes given a DEM, without requiring an underlying circuit noise model. Next, we demonstrate the utility of DEM estimation in discovering and investigating errors not explained by typical circuit noise. Finally, we depart from the direct discussion of DEMs to present large, anomalous correlations in the Google data.
@@ I.1 Prior Work
We now present a brief, incomplete chronology of the concept of detector error models and error estimation using syndrome data. We formally define detectors 2024DerksDesigning in the context of QEC below, but informally, a detector is a parity constraint among measurement outcomes that is deterministic in the absence of noise. A given class of (detectable) errors violates, or flips, at least one parity constraint, and a DEM specifies a) the set of detectors flipped by each error class—the structure—and b) the independent rate of each error class—the parameters.
The DEM estimation problem was first directly addressed for a DEM with graphical structure (in which each error class flips at most a pair of detectors) in https://doi.org/10.1002/qute.201800012, constructing the analytically estimable $p_{ij}$ which appear in later works. This work is extended in PhysRevLett.128.110504, which includes a concise description of the DEM in terms of a decoding hypergraph in Section II and describes an algorithm for estimation of error parameters of a fixed hypergraph in Supplement G. These approaches to the problem of rate estimation are the conceptual ancestors of one of the two classes of algorithms discussed here: those based on moments of detectors.
Meanwhile, the DEM was identified as a first class object for stabilizer simulation and QEC decoding via minimum-weight perfect matching in gidney2021stim and Higgot_2025 respectively. We note that the definitions adopted in these works are the ones most closely followed in our work due primarily to use of the tool chain described therein.
In PhysRevResearch.3.013292, the authors discuss many of the ideas covered here and present a Pauli error-process estimator utilizing a modified belief-propagation (BP) decoder and expectation-maximization (EM) algorithm. However, in that and a following work Wagner2022paulichannelscanbe, they limit their approach to the “framework of quantum-data syndrome codes” where Pauli errors may occur only on data qubits and syndrome measurements may err. This restriction is not observed in our work.
@@
Two recent works mark an important advance in understanding DEMs: the analysis of detector parities, rather than moments, for DEM learning. Remm et al. remm2025 introduced a parity-based algorithm for rate estimation and applied it to estimate a DEM from syndromes generated by a laboratory QEC chip. By estimating rates for a DEM with hand-picked structure, the authors diagnose leakage and multi-qubit errors in the chip. The recent work of blumekohout2025estimatingdetectorerrormodels, independent of remm2025, made great theoretical strides in relating parities of sets of detectors to DEM parameters via a Walsh-Hadamard transform, thereby linking DEM research to a rich body of literature and laying the groundwork for general algorithms for both rate estimation and structure learning.
@@ II Detector Error Models and some Mathematical Properties
We are indebted to 2024DerksDesigning; blumekohout2025estimatingdetectorerrormodels; gidney2021stim; Higgot_2025 from which we liberally borrow. We remind the reader that a detector is a parity constraint on a series of measurement outcomes in a QEC circuit. For the repetition and surface codes discussed here, a detector is typically the mod-2 sum of the same stabilizer measurement in successive rounds of syndrome extraction. Thus, in the time-bulk of a memory experiment (i.e. except for the initial and final rounds), a detector exists for each stabilizer in each round and takes the value one if and only if the corresponding stabilizer measurement changed in that round. We dispense with raw stabilizer measurements and define a syndrome as the sequence of detector values for one replicate, or shot, of a multi-round, logical memory experiment.
@@
Given $(\mathbf{M},\bm{\theta})$, we generate data by drawing a random excitation vector $\mathbf{e}$ with $e_{i}\sim$ Bernoulli($\theta_{i}$). Then $\mathbf{x}=\mathbf{Me},$ where the matrix-vector multiply is performed in $\mathbb{F}_{2}$. This is essentially the mathematical formulation for the process of QEC simulation in stim gidney2021stim. Linear algebra in $\mathbb{F}_{2}$ enforces the English language statement, “the probability a given detector flips is the probability an odd number of excited hyperedges touch a detector.” Importantly, in this formulation the elements of $\mathbf{e}$ are all independent.
@@ II.1 Moments, Parities, Polarizations and More
$\mu_{\mathbf{y}}\equiv\langle[\mathbf{y}\leq\mathbf{x}]\rangle=\sum_{\mathbf{x}\in\mathbb{F}_{2}^{n}}[\mathbf{y}\leq\mathbf{x}]p_{\mathbf{x}}=\sum_{\mathbf{e}\in\mathbb{F}_{2}^{E}}[\mathbf{y}\leq\mathbf{M}\mathbf{e}]=\Pr(\mathbf{e}|\bm{\theta})=\sum_{\mathbf{e}\in\mathbb{F}_{2}^{E}}[\mathbf{y}\leq\mathbf{M}\mathbf{e}]\prod_{i\in[E]}\theta_{i}^{e_{i}}(1-\theta_{i})^{1\oplus e_{i}}.$ (5)
@@
This is equivalent to Equation 5 in PhysRevLett.128.110504 from which they derive their estimation algorithm. These equations also form the basis for the recent work takou2025estimatingdecodinggraphshypergraphs, however the authors do not appear to have directly addressed (nor significantly suffered from) the issue of potentially many excitations now having the same signature within a moment. We derive a distinct estimation algorithm from both approaches using Equation 6.
Following blumekohout2025estimatingdetectorerrormodels we examine the parity of a subset of detectors indicated by $\mathbf{y}$, but under the group isomorphism $\{\{0,1\},\oplus\}\to\{\{+1,-1\},\times\}$,
@@
As in blumekohout2025estimatingdetectorerrormodels, we define the depolarization,
@@
This theorem has two interesting consequences. First, it provides a derivation of Equation 10 of remm2025, which the authors verify but do not derive. Exponentiation of Equation 17 yields,
@@
Equation 25 is equivalent to Equation 31 in blumekohout2025estimatingdetectorerrormodels accounting for a different normalization convention and variable names.
We note that the LHS of Equation 24 and RHS of Equation 16 are not equal for arbitrary $\bm{\psi}$. The authors of blumekohout2025estimatingdetectorerrormodels note that $\mathbf{W}$ is not invertible, as it annihilates the vector $(1,0,0,\cdots,0)$. This leaves $\psi_{0}$ as a free parameter. Applying Equation 17 with $Y=\emptyset$ yields
@@
Using this value of $\psi_{0}$, Equation 24 and Equation 16 yield equivalent $\bm{\omega}$. The authors of blumekohout2025estimatingdetectorerrormodels define the negative of this quantity as the total attenuation. By taking this value for an otherwise unused parameter, we exchange the non-invertible binary Walsh transform for the traditional Hadamard transform and obtain a bijection from attenuations to depolarizations.
@@ II.3 DEM Constraints
shows $0\leq f(\theta_{01},\theta_{10})\leq 1$. Thus $0\leq\mathsf{Cov}\leq 1/4$. Therefore, in a two-bit DEM, small errors cannot produce anti-correlations between the two detectors. In Section 3.1 of blumekohout2025estimatingdetectorerrormodels, the authors show that the covariance between detectors $i$ and $j$ is proportional to the aggregated attenuation of event $\{i,j\}$, which is in turn the sum of attenuations for all DEM events for which $i,j=1$;
@@
The entries $\frac{1}{2}-\frac{1}{2}\sqrt{1-\frac{4}{3}p}$ are identical to the value produced as a solution to independent error processes representing uniform depolarizing single qubit errors in Gidney2020decorrelated-depolarization and used widely in stim.
@@ II.4 Further Reading
The well studied class of log-linear models for multivariate binary distributions is defined by wasserman2003statistics to satisfy:
@@
Note that in log-linear models $\lambda_{A}=0$, if and only if there exists a partition of $A=\{A_{1},A_{2}\}$ such that bits in $A_{1}$ are conditionally independent of bits in $A_{2}$ given the bits in [n]A[n]-A. Additionally, for a subclass of such models—chordal, graphical models—there is an efficient algorithm for structure learning Petitjean2015-SDM.
The class of log-mean linear models introduced in 2013_log-mean_linear_models_for_binary_data is a parameterization of discrete binary variable distributions defined by the parameter vector
@@
However, hendy1993 chose to call it the Hadamard conjugation transform when using this to perform a similar estimation problem on phylogenetic trees based on observed gene frequencies. Unlike the other two models above, we must account for the pole when $\mathbf{Hp}$ has a component equal to zero. While the authors know of no general condition to avoid this, when $p_{0}>\sum_{i=1}^{2^{n}-1}p_{i}$ and all $p_{i}$ are non-zero there are no divergences. Whether this presents a hurdle to (non-physical) log-polarization linear models providing a complete description of binary string distributions remains to be seen. However, there is a very clear interpretation of the meaning of zeros in $\bm{\psi}$ (at least in the physical case); there is no process which flips exactly the bits indicated by that component.
@@ III.1 Moment-Based Algorithms
and solving using a root-finder, such as the options available in scipy.optimize 2020SciPy-NMeth.
@@
We also may wish to learn the structure of $\mathbf{M}$. We present Algorithm 2, a moment-based algorithm for doing this in a limited case of low maximal hyperedge cardinality. To begin, this algorithm finds statistically significant correlations from among all pairs of detectors using the analysis introduced in https://doi.org/10.1002/qute.201800012 and popularized in google2021exponentialsuppression. Here, we reproduce Equations 11 and 19 from the Supplementary Material of google2021exponentialsuppression in our notation with minor algebraic changes:
@@ III.2 Parity Based Algorithms
Two works remm2025; blumekohout2025estimatingdetectorerrormodels have recently suggested algorithms for estimating DEMs from syndromes via parities. While remm2025 does not make their estimation algorithm explicit, we believe it is equivalent to that presented in Section 4.1 of blumekohout2025estimatingdetectorerrormodels. We demonstrate our reasoning below before providing pseudo-code. Return to Equation 17, but now assume that $S$ has the highest cardinality of any hyperedge in the DEM, i.e. $S\subset A\implies\psi_{A}=0$, resulting in
@@
using attenuations from previous iterations to compute the second summation in the RHS. A logarithm and some manipulation changes this to exactly Equation 10 from remm2025 as shown in (21).
@@
To perform structure learning, we present Algorithm 4, noting the same requirement for maximal hyperedge cardinality to be relatively low. This algorithm is structurally similar to Algorithm 2. Both structure learning algorithms are likely to benefit from better hypothesis testing to prune hyperedges as in 10.1145/2939672.2932775.
@@
We also examine the possibility of rate estimation from parities from pseudo-inversion of a tall matrix a.k.a least-squares regression as suggested in blumekohout2025estimatingdetectorerrormodels. The benefit of this approach is an algorithm which is no longer exponential in the weight of hyperedges in the DEM. In particular we must take at least $E$ distinct samples to reach a full-rank design matrix, after which the computational complexity is determined by the cost of least-squares estimation or pseudo-inverse calculation. In some cases, this can be significantly faster than Algorithm 3. We note in the next section some subtleties that arise due to the noise experienced by depolarizations, which prevent these algorithms from working for all DEMs and place non-obvious constraints on the set of depolarizations one may wish to calculate.
@@ III.5 DEM Estimate Variances
The final variance estimation method uses the jackknife wu1986jacknife. For each of $R$ replicates, we perform Algorithm 3 on $\mathbf{X}$ with a single, randomly chosen shot omitted, resulting in a set of estimates for each hyperedge excitation rate: $\{\hat{\theta}_{\mathbf{s}}[i]\}_{i=1}^{R}$. The variance may then be estimated as
@@
In a traditional jackknife jackknife each sample is omitted exactly once, yielding a deterministic estimate of variance with $R=N$. However, with syndrome corpora routinely exceeding one million shots, the traditional jackknife would be prohibitively expensive. Since it is competing with methods that are already quite fast, we choose $R=10^{3}$, which requires about an order of magnitude more computation than estimating the excitation rates themselves. Even with this setting, the jackknife is both slower and less accurate than the other two methods.

### v2

@@ I Introduction
Over the last year, teams have published results directly demonstrating improved performance of logically encoded circuits Reichardt et al. (2024) and increased lifetimes of error corrected states AI and Collaborators (2024). This leads naturally to an increased interest in the analysis of the resulting syndrome data and the description of the data generating process via detector error models (DEMs).
@@
We then provide algorithms for two learning tasks: a) rate estimation, in which the structure of the DEM is given and the goal is to learn the parameters from syndrome data-sets, and b) structure learning, where both structure and parameters must be learned. For each task, we describe two algorithmic approaches, one based on moments of detectors and the other based on parities of the same. In the last section, we demonstrate the application of these algorithms to recent Google experiments AI and Collaborators (2024) and on simulated data. Specifically, we discuss the accuracy of decoding using estimated priors versus non-informative priors and priors trained to minimize the logical error rate (LER). Then, we evaluate the agreement between various DEMs and syndromes using a method for estimating the likelihood of syndromes given a DEM, without requiring an underlying circuit noise model. Next, we demonstrate the utility of DEM estimation in discovering and investigating errors not explained by typical circuit noise. Finally, we depart from the direct discussion of DEMs to present large, anomalous correlations in the Google data.
@@ I.1 Prior Work
We now present a brief, incomplete chronology of the concept of detector error models and error estimation using syndrome data. We formally define detectors et al (2024) in the context of QEC below, but informally, a detector is a parity constraint among measurement outcomes that is deterministic in the absence of noise. A given class of (detectable) errors violates, or flips, at least one parity constraint, and a DEM specifies a) the set of detectors flipped by each error class—the structure—and b) the independent rate of each error class—the parameters.
The DEM estimation problem was first directly addressed for a DEM with graphical structure (in which each error class flips at most a pair of detectors) in Spitz et al. (2018), constructing the analytically estimable $p_{ij}$ which appear in later works. This work is extended in Chen et al. (2022), which includes a concise description of the DEM in terms of a decoding hypergraph in Section II and describes an algorithm for estimation of error parameters of a fixed hypergraph in Supplement G. These approaches to the problem of rate estimation are the conceptual ancestors of one of the two classes of algorithms discussed here: those based on moments of detectors.
Meanwhile, the DEM was identified as a first class object for stabilizer simulation and QEC decoding via minimum-weight perfect matching in Gidney (2021) and Higgott and Gidney (2025) respectively. We note that the definitions adopted in these works are the ones most closely followed in our work due primarily to use of the tool chain described therein.
In Wagner et al. (2021), the authors discuss many of the ideas covered here and present a Pauli error-process estimator utilizing a modified belief-propagation (BP) decoder and expectation-maximization (EM) algorithm. However, in that and a following work Wagner et al. (2022), they limit their approach to the “framework of quantum-data syndrome codes” where Pauli errors may occur only on data qubits and syndrome measurements may err. This restriction is not observed in our work.
@@
Two recent works mark an important advance in understanding DEMs: the analysis of detector parities, rather than moments, for DEM learning. Remm et al. Remm et al. (2025) introduced a parity-based algorithm for rate estimation and applied it to estimate a DEM from syndromes generated by a laboratory QEC chip. By estimating rates for a DEM with hand-picked structure, the authors diagnose leakage and multi-qubit errors in the chip. The recent work of Blume-Kohout and Young (2025), independent of Remm et al. (2025), made great theoretical strides in relating parities of sets of detectors to DEM parameters via a Walsh-Hadamard transform, thereby linking DEM research to a rich body of literature and laying the groundwork for general algorithms for both rate estimation and structure learning.
@@ II Detector Error Models and some Mathematical Properties
We are indebted to et al (2024); Blume-Kohout and Young (2025); Gidney (2021); Higgott and Gidney (2025) from which we liberally borrow. We remind the reader that a detector is a parity constraint on a series of measurement outcomes in a QEC circuit. For the repetition and surface codes discussed here, a detector is typically the mod-2 sum of the same stabilizer measurement in successive rounds of syndrome extraction. Thus, in the time-bulk of a memory experiment (i.e. except for the initial and final rounds), a detector exists for each stabilizer in each round and takes the value one if and only if the corresponding stabilizer measurement changed in that round. We dispense with raw stabilizer measurements and define a syndrome as the sequence of detector values for one replicate, or shot, of a multi-round, logical memory experiment.
@@
Given $(\mathbf{M},\bm{\theta})$, we generate data by drawing a random excitation vector $\mathbf{e}$ with $e_{i}\sim$ Bernoulli($\theta_{i}$). Then $\mathbf{x}=\mathbf{Me},$ where the matrix-vector multiply is performed in $\mathbb{F}_{2}$. This is essentially the mathematical formulation for the process of QEC simulation in stim Gidney (2021). Linear algebra in $\mathbb{F}_{2}$ enforces the English language statement, “the probability a given detector flips is the probability an odd number of excited hyperedges touch a detector.” Importantly, in this formulation the elements of $\mathbf{e}$ are all independent.
@@ II.1 Moments, Parities, Polarizations and More
$\mu_{\mathbf{y}}\equiv\langle[\mathbf{y}\leq\mathbf{x}]\rangle=\sum_{\mathbf{x}\in\mathbb{F}_{2}^{n}}[\mathbf{y}\leq\mathbf{x}]p_{\mathbf{x}}=\sum_{\mathbf{e}\in\mathbb{F}_{2}^{E}}[\mathbf{y}\leq\mathbf{M}\mathbf{e}]\Pr(\mathbf{e}|\bm{\theta})=\sum_{\mathbf{e}\in\mathbb{F}_{2}^{E}}[\mathbf{y}\leq\mathbf{M}\mathbf{e}]\prod_{i\in[E]}\theta_{i}^{e_{i}}(1-\theta_{i})^{1\oplus e_{i}}.$ (5)
@@
This is equivalent to Equation 5 in Chen et al. (2022) from which they derive their estimation algorithm. These equations also form the basis for the recent work Takou and Brown (2025), however the authors do not appear to have directly addressed (nor significantly suffered from) the issue of potentially many excitations now having the same signature within a moment. We derive a distinct estimation algorithm from both approaches using Equation 6.
Following Blume-Kohout and Young (2025) we examine the parity of a subset of detectors indicated by $\mathbf{y}$, but under the group isomorphism $\{\{0,1\},\oplus\}\to\{\{+1,-1\},\times\}$,
@@
As in Blume-Kohout and Young (2025), we define the depolarization,
@@
This theorem has two interesting consequences. First, it provides a derivation of Equation 10 of Remm et al. (2025), which the authors verify but do not derive. Exponentiation of Equation 17 yields,
@@
Equation 25 is equivalent to Equation 31 in Blume-Kohout and Young (2025) accounting for a different normalization convention and variable names.
We note that the LHS of Equation 24 and RHS of Equation 16 are not equal for arbitrary $\bm{\psi}$. The authors of Blume-Kohout and Young (2025) note that $\mathbf{W}$ is not invertible, as it annihilates the vector $(1,0,0,\cdots,0)$. This leaves $\psi_{0}$ as a free parameter. Applying Equation 17 with $Y=\emptyset$ yields
@@
Using this value of $\psi_{0}$, Equation 24 and Equation 16 yield equivalent $\bm{\omega}$. The authors of Blume-Kohout and Young (2025) define the negative of this quantity as the total attenuation. By taking this value for an otherwise unused parameter, we exchange the non-invertible binary Walsh transform for the traditional Hadamard transform and obtain a bijection from attenuations to depolarizations.
@@ II.3 DEM Constraints
shows $0\leq f(\theta_{01},\theta_{10})\leq 1$. Thus $0\leq\mathsf{Cov}\leq 1/4$. Therefore, in a two-bit DEM, small errors cannot produce anti-correlations between the two detectors. In Section 3.1 of Blume-Kohout and Young (2025), the authors show that the covariance between detectors $i$ and $j$ is proportional to the aggregated attenuation of event $\{i,j\}$, which is in turn the sum of attenuations for all DEM events for which $i,j=1$;
@@
The entries $\frac{1}{2}-\frac{1}{2}\sqrt{1-\frac{4}{3}p}$ are identical to the value produced as a solution to independent error processes representing uniform depolarizing single qubit errors in Gidney (2020) and used widely in stim.
@@ II.4 Further Reading
The well studied class of log-linear models for multivariate binary distributions is defined by Wasserman (2003) to satisfy:
@@
Note that in log-linear models $\lambda_{A}=0$, if and only if there exists a partition of $A=\{A_{1},A_{2}\}$ such that bits in $A_{1}$ are conditionally independent of bits in $A_{2}$ given the bits in [n]A[n]-A. Additionally, for a subclass of such models—chordal, graphical models—there is an efficient algorithm for structure learning Petitjean and Webb (2015).
The class of log-mean linear models introduced in Roverato et al. (2013) is a parameterization of discrete binary variable distributions defined by the parameter vector
@@
However, Hendy and Penny (1993) chose to call it the Hadamard conjugation transform when using this to perform a similar estimation problem on phylogenetic trees based on observed gene frequencies. Unlike the other two models above, we must account for the pole when $\mathbf{Hp}$ has a component equal to zero. While the authors know of no general condition to avoid this, when $p_{0}>\sum_{i=1}^{2^{n}-1}p_{i}$ and all $p_{i}$ are non-zero there are no divergences. Whether this presents a hurdle to (non-physical) log-polarization linear models providing a complete description of binary string distributions remains to be seen. However, there is a very clear interpretation of the meaning of zeros in $\bm{\psi}$ (at least in the physical case); there is no process which flips exactly the bits indicated by that component.
@@ III.1 Moment-Based Algorithms
and solving using a root-finder, such as the options available in scipy.optimize Virtanen et al. (2020).
@@
We also may wish to learn the structure of $\mathbf{M}$. We present Algorithm 2, a moment-based algorithm for doing this in a limited case of low maximal hyperedge cardinality. To begin, this algorithm finds statistically significant correlations from among all pairs of detectors using the analysis introduced in Spitz et al. (2018) and popularized in 5. Here, we reproduce Equations 11 and 19 from the Supplementary Material of 5 in our notation with minor algebraic changes:
@@ III.2 Parity Based Algorithms
Two works Remm et al. (2025); Blume-Kohout and Young (2025) have recently suggested algorithms for estimating DEMs from syndromes via parities. While Remm et al. (2025) does not make their estimation algorithm explicit, we believe it is equivalent to that presented in Section 4.1 of Blume-Kohout and Young (2025). We demonstrate our reasoning below before providing pseudo-code. Return to Equation 17, but now assume that $S$ has the highest cardinality of any hyperedge in the DEM, i.e. $S\subset A\implies\psi_{A}=0$, resulting in
@@
using attenuations from previous iterations to compute the second summation in the RHS. A logarithm and some manipulation changes this to exactly Equation 10 from Remm et al. (2025) as shown in (21).
@@
To perform structure learning, we present Algorithm 4, noting the same requirement for maximal hyperedge cardinality to be relatively low. This algorithm is structurally similar to Algorithm 2. Both structure learning algorithms are likely to benefit from better hypothesis testing to prune hyperedges as in Webb and Petitjean (2016).
@@
We also examine the possibility of rate estimation from parities from pseudo-inversion of a tall matrix a.k.a least-squares regression as suggested in Blume-Kohout and Young (2025). The benefit of this approach is an algorithm which is no longer exponential in the weight of hyperedges in the DEM. In particular we must take at least $E$ distinct samples to reach a full-rank design matrix, after which the computational complexity is determined by the cost of least-squares estimation or pseudo-inverse calculation. In some cases, this can be significantly faster than Algorithm 3. We note in the next section some subtleties that arise due to the noise experienced by depolarizations, which prevent these algorithms from working for all DEMs and place non-obvious constraints on the set of depolarizations one may wish to calculate.
@@ III.5 DEM Estimate Variances
The final variance estimation method uses the jackknife Wu (1986). For each of $R$ replicates, we perform Algorithm 3 on $\mathbf{X}$ with a single, randomly chosen shot omitted, resulting in a set of estimates for each hyperedge excitation rate: $\{\hat{\theta}_{\mathbf{s}}[i]\}_{i=1}^{R}$. The variance may then be estimated as
@@
In a traditional jackknife Tukey (1958) each sample is omitted exactly once, yielding a deterministic estimate of variance with $R=N$. However, with syndrome corpora routinely exceeding one million shots, the traditional jackknife would be prohibitively expensive. Since it is competing with methods that are already quite fast, we choose $R=10^{3}$, which requires about an order of magnitude more computation than estimating the excitation rates themselves. Even with this setting, the jackknife is both slower and less accurate than the other two methods.